# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ 4 Implementation and Experiments
We have implemented a tool called TupInf for STP inference based on the algorithm in Section 2, and a new CHC solver called CHoCoL based on the method in Section 3, using TupInf as a backend. Section 4.1 gives an overview of those tools, and Section 4.2 reports experimental results.
@@ 4.1.1 TupInf: a Solvable Tuple Pattern Inference Tool
TupInf has been implemented in OCaml. TupInf takes learning data in the csv format, and outputs tuple patterns as defined in Section 2.1 with the extensions described in Sections 2.3.1 (postfix), 2.3.2 (reverse), and 2.3.4 (set and multiset patterns). The alphabet $\Sigma$ used in the learning data consists of digits and upper/lower characters of the English alphabet. There is an option to disable/enable constant patterns (i.e., R-CPrefix and R-CPostfix); it was disabled in the experiments reported below. A web demo interface of TupInf is available at https://www.kb.is.s.u-tokyo.ac.jp/~koba/tupinf/.
@@ 4.1.2 CHoCoL: a CHC Solver for List-like Data Structures
CHoCoL (Constrained Horn Clauses over Lists) is also implemented in OCaml and uses TupInf for STP inference, Z3 and CVC5 for SMT solving, and HoIce [8] for CHC solving over integer arithmetic, as backend solvers. Currently, CHoCoL supports only list-like data structures as ADTs. More precisely, it supports ADTs that have two constructors: one with no arguments and the other with two arguments whose types are some element type and the ADT itself. CHoCoL also supports natural numbers as lists with unit elements.
CHoCoL has four internal modes: list-mode, set-mode, multiset-mode, and list-len-mode. In the list-mode, it infers tuple patterns for list-like data structures as described in Section 3.2. In the list-len-mode, it additionally infers invariants on the lengths of lists as described in Section 3.3 by using the backend CHC solver. Then, CHoCoL checks whether the obtained invariants are strong enough to imply the goal clause, by using the backend SMT solver. Besides the support of multiple predicate variables, the procedure in the list-mode deviates from Procedure 1 in Section 3 in the following points: (1) We skip line 5 to deduce unsatisfiability; the unsatisfiability is checked by a separate procedure as described below. (2) We do not accumulate the tuple patterns found so far (cf. lines 8–9); so $T$ is a singleton set $\{t_{\mathit{new}}\}$ on line 10, and if $T\models G$ does not hold on line 10, we just proceed to the next step for inferring invariants on list lengths, to strengthen the invariants represented by $t_{\mathit{new}}$. (3) $\mathtt{STPinf}{}$ on line 7 is actually deterministic. (4) To collect more samples on line 13, we use only the counterexample-guided method, which does not satisfy the second assumption in Theorem 3.3 (recall Remark 3.2). Due to these differences, the implementation does not satisfy the relative completeness as stated in Theorem 3.3. That does not, however, seem to be a significant issue in practice, according to the experimental results in Section 4.2.
In the set-mode (resp. multiset-mode), CHoCoL infers set (resp. multiset) tuple patterns as described in Section 3.4 and checks whether inferred invariants imply the goal clause. Given a system of CHCs as input, CHoCoL tries the four modes (list-mode, set-mode, multiset-mode, and list-len-mode in this order) one by one until it finds that the CHCs are satisfiable.
Notice that the procedure explained so far can only prove the satisfiability of the CHCs. CHoCoL also runs in parallel a procedure based on symbolic execution to prove the unsatisfiability. Given a CHC whose goal clause is $\bigwedge_{i}P_{i}(\widetilde{y}_{i})\land C\Rightarrow\mathit{false}$, CHoCoL searches for an assignment to variables $\widetilde{y}_{i}$ that makes the body ($\bigwedge_{i}P_{i}(\widetilde{y}_{i})\land C$) of the goal clause valid by recursively unfolding the predicates in the goal in a breadth-first manner. The details of this procedure are omitted, since it is rather straightforward and contains few novel ideas. It would be possible to utilize the result of tuple pattern inference also to speed-up the refutation process, but that is left for future work.
@@ 4.2 Experiments
To evaluate the effectiveness of TupInf and CHoCoL, we have conducted experiments on a benchmark set selected from CHC-COMP 2025 ADT-LIN category. [^10] https://chc-comp.github.io/ The benchmark set consists of 445 instances of the CHC satisfiability problem, which are all the instances of the ADT-LIN category where the used ADTs are only list-like data structures. The experiments were conducted on a machine with AMD Ryzen 9 5900X CPU and 32GB RAM. The cpu and wallclock time limits were set to 3 minutes, and the memory limit was set to 8GB. The number of CPU cores was limited to 4.
We compared CHoCoL with the state-of-the-art CHC solvers that support ADTs: RInGen [33], Spacer/Racer [32, 28], Eldarica [24], HoIce [8, 9], and Catalia [30]. The results are shown in Table 1 and Figure 1. To confirm that our method serves as a complement to previous methods, we have also prepared a combination of CHoCoL and Catalia, shown as CHoCoL+Catalia in the table and figure; it runs CHoCoL and Catalia in parallel. [^11] A combination of CHoCoL and Catalia, called “ChocoCatalia” won the ADT-LIN category of CHC-COMP 2025; the configuration of CHoCoL+Catalia reported in this submission slightly differs from “ChocoCatalia”.
@@
CHoCoL 167 (84) 80 (0) 247 (84)
RInGen 64 (23) 37 (4) 101 (27)
Spacer 20 (2) 90 (0) 110 (2)
Eldarica 20 (0) 84 (0) 104 (0)
HoIce 24 (6) 43 (0) 67 (6)
Catalia 87 (12) 87 (0) 174 (12)
CHoCoL+Catalia 188 84 272
@@
Table 1 shows the numbers of solved instances by CHoCoL and the other CHC solvers. For each solver, the column Solved (SAT) (resp. Solved (UNSAT)) shows the number of satisfiable (resp. unsatisfiable) instances solved, and the column Solved (all) shows the total number of instances solved. The numbers in parentheses are the numbers of uniquely solved instances, i.e., instances that the solver could solve but no other solvers (except for CHoCoL+Catalia) could solve within the time limit.
The experimental results demonstrate the effectiveness of CHoCoL. CHoCoL achieved the highest total number of solved instances (247) among all compared solvers, showing its overall superiority in handling CHCs with list-like data structures. A key strength of CHoCoL is its performance on satisfiable instances, where it solved 167 instances compared to 87 by Catalia and 64 by RInGen. More importantly, CHoCoL uniquely solved 84 satisfiable instances that no other solver could handle within the time limit,121212 A increase of the time limit would not significantly change the situation, because the benchmark set contains a number of instances that require reasoning about list equalities, as in the examples in Section 1, which cannot be solved by Catalia and RInGen, which abstract lists to integers or elements of a finite domain. demonstrating the effectiveness of the STP inference approach for CHC solving.
The result of CHoCoL+Catalia, though, shows that CHoCoL also has a limitation, and that it is better to use it as a complement to other solvers. Also, in contrast with the strong performance of CHoCoL for SAT instances, CHoCoL showed slightly lower performance than other solvers (except RInGen) for UNSAT instances. This is unsurprising because our current refutation procedure is quite naive, which does not use the result of the STP inference at all. As mentioned in Section 4.1, we plan to apply SPT inference also to improve the refutation procedure.
@@
Total 167 84
Table 2 shows the numbers of solved SAT instances by each mode of CHoCoL. The row “refutation” shows the number of instances found to be satisfiable only by finite unfolding. It is remarkable that the list-mode alone solved more SAT instances than the other solvers. As shown in the row “list-len-mode”, the combination of STPs and CHC solving over integer arithmetic was also quite effective.
We have also manually inspected some of the SAT instances that could not be solved by CHoCoL. They typically involve functions such as sort, filter, map, and fold, whose input/output relations cannot be expressed by STPs. An extension with the map relation is briefly discussed in Appendix B, but it has not yet been implemented, and an extension with the sort relation loses desirable properties of STPs, as discussed in Section 2.3.3. Handling these features is left for future work.
@@ 5.1 Learning Languages from Positive Samples
The problem of learning a language from only positive samples [18] has been actively studied, and various language classes have been found to be learnable. As already mentioned, most closely related to our notion of tuple patterns is Angluin’s pattern languages [2, 1] and its variations [49, 50]. Indeed, “erasing” pattern languages [49, 50] (in which empty strings may be substituted for variables, as opposed to Angluin’s original pattern languages) can be considered singleton tuple patterns, and conversely, if we introduce a special symbol \$ (which must not occur in strings substituted for variables; such a restriction may be expressed by patterns with regular constraints [49]), then a tuple pattern $(p_{1},\ldots,p_{n})$ can be expressed as the pattern $p_{1}\$\cdots\$p_{n}$. No polynomial time learning algorithm is known for the full class of pattern languages (and in fact, the existence of such an algorithm is highly unlikely given the result of  [1]). Various subclasses of pattern languages are known to be polynomial time learnable from positive data [1, 48, 44, 49, 39]. Among others, Angluin [1] and Shinohara [50] gave polynomial time algorithms respectively for patterns consisting of a single variable and for “regular” (erasing) patterns where each variable may occur only once. Mitchell [39] extended the latter result to quasi-regular patterns, where variables must occur the same number of times. Those subclasses are not large enough to express tuple patterns like $(xy,xyx)$ (which contains two variables, each of which occurs a different number of times).
@@ 5.2 Data-Driven Approaches to Program Verification
Various data-driven approaches to program verification have recently been proposed, especially for automated discovery of program invariants. Some of them require only positive samples [47, 25], some require negative samples as well [61, 60, 46], and some require even implication examples of the form $d_{1}\land\cdots\land d_{k}\Rightarrow d$, which means “if $d_{1},\ldots,d_{k}$ are positive samples, then $d$ is also positive” [17, 8, 16, 31]. The used techniques vary from algebraic ones [47, 25], SVM [61, 60], and neural networks [46, 31]. Those techniques have been used mainly for the discovery of invariants on integers, and it is unclear how they can be adopted to discover invariants on list-like data structures.
Among them, the techniques of Sharma et al. [47] and Ikeda et al. [25] are technically closest to ours (although they are for discovering numerical equality constraints), in that those methods can also quickly infer equality constraints only from a small number of examples. In fact, our STP inference algorithm, which repeatedly simplifies learning data in matrix form, has been inspired by those techniques. Ikeda et al. [25] applied that technique to CHC solving. The application to CHC solving is less trivial in our context, however. In the case of Ikeda et al. [25], which finds equality constraints on integers, it is sufficient to insert those invariants as a part of preprocessing, and then invoke an off-the-shelf CHC solver. For example, if we can find an invariant $P(x,y)\Rightarrow x+2y=3$, then we can provide that information to a CHC solver, just by replacing every occurrence of $P(z,w)$ in CHC with $P(z,w)\land z+2w=3$. In contrast, since there are no CHC solvers that can directly reason about sequence equality constraints, we had to develop a more sophisticated method for applying the result of STP inference to CHC solving, as discussed in Section 3.
@@ 5.3 CHC Solving for Algebraic Data Types
Solving CHCs over algebraic data types (or, automated verification of programs manipulating ADTs) often require finding and reasoning about inductive properties. To this end, Unno et al. [54] introduced an automated method for applying induction in CHC solving. De Angelis et al. [14] later proposed a method to achieve similar effects without explicit induction, by using unfold/fold transformations. Mordvinov and Fedykovich [40] also proposed a variation of unfold/fold transformations. Losekoot et al. [34, 35] applied automata-based techniques to reason about relational properties of ADTs, which also have a similar effect to unfold/fold transformation-based techniques. These methods tend to be fragile, requiring a lot of heuristics to automatically decide how and when to apply induction, unfold/fold transformations, etc.
Several methods for CHC solving have also been proposed, which abstract ADTs to elements of a finite domain or integers. Kostyukov [33] proposed a method for abstracting ADTs using automata, while the Eldarica CHC solver [24] abstracts ADTs by their size. Govind V. K. et al. [28] proposed a method that uses recursively defined functions (RDTs) over ADTs for abstraction, but those RDTs must be explicitly specified. Katsura et al. [30] proposed a method for automatically finding a catamorphism-based abstraction for ADTs. While these methods are effective for certain CHCs, they are not effective for proving the equality of ADTs; although the abstraction by using an abstraction map $\alpha$ may be used for proving $\alpha(l_{1})=\alpha(l_{2})$, it does not imply $l_{1}=l_{2}$ in general. Thus, those methods cannot be used to prove neither of the examples in Section 1. As confirmed in our experiments, our method using STPs is complementary to those methods; in fact, the combination of our method with Catalia achieved the best performance in the experiments in Section 4, and also in the competition CHC-COMP 2025.
@@ 6 Conclusions
We have introduced a new class of languages, called solvable tuple patterns,131313 A (formal) “language” usually refers to a set of strings, while a solvable tuple pattern expresses a set of tuples consisting of strings; note, however, a tuple $(s_{1},\ldots,s_{n})$ may be viewed as a string $s_{1}\$\cdots\$s_{n}$. that are efficiently learnable from only positive samples, proposed an inference algorithm, and proved that SPTs and the inference algorithm satisfy attractive properties. We have applied STPs to automatic verification of programs manipulating list-like data structures. We have implemented a new CHC solver called CHoCoL based on the proposed approach, and confirmed the effectiveness of the approach through experiments.
Future work includes: (i) extension of STPs to deal with tree-like data structures, (ii) extension of STPs to infer Boolean combinations of STPs, like $(x,y,xz)\lor(x,y,yz)$ (which means “either the first or second element is a prefix of the third element), (iii) applications of other classes of languages learnable from positive samples to CHC solving, and (iii) extension and generalization of the method to combine CHC solvers over different domains.

### v2

@@ 6 Implementation and Experiments
We have implemented a tool called TupInf for (C)STP inference based on the algorithm in Section 3, and a new CHC solver called CHoCoL based on the method in Section 5, using TupInf as a backend. Section 6.1 gives an overview of those tools, and Section 6.2 reports experimental results.
@@ 6.1.1 TupInf: a Solvable Tuple Pattern Inference Tool
TupInf has been implemented in OCaml. TupInf takes learning data in the csv format, and outputs STPs or CSTPs with the extensions described in Sections 4.1 (reverse), and 4.2 (set and multiset patterns). The alphabet $\Sigma$ used in the learning data consists of digits and uppercase and lowercase English letters. TupInf provides options to (i) enable or disable constant patterns (i.e., R-CPrefix and R-CSuffix), (ii) choose between STPs and CSTPs, and (iii) select between sequence, multiset, and set patterns. The option for constant patterns was disabled in the experiments reported below.
@@ 6.1.2 CHoCoL: a CHC Solver for List-like Data Structures
CHoCoL (Constrained Horn Clauses over Lists) is also implemented in OCaml and uses TupInf for CSTP inference, Z3 and CVC5 for SMT solving, and HoIce [7] for CHC solving over integer arithmetic, as backend solvers. Currently, CHoCoL supports only list-like data structures as ADTs. More precisely, it supports ADTs that have two constructors: one with no arguments and the other with two arguments whose types are some element type and the ADT itself. CHoCoL also supports natural numbers represented as lists with unit elements. CHoCoL has five internal modes: list-stp-mode, set-mode, multiset-mode, list-len-mode, and list-cstp-mode. Given a system of CHCs as input, CHoCoL tries five modes (in the order listed above) one by one until it finds that the CHCs are satisfiable. In the list-stp-mode and list-cstp-mode, the tool infers tuple patterns for list-like data structures as described in Section 5.1, but the inferred patterns are restricted to STPs in the former for the sake of efficiency. (Experimental results suggest that CSTP inference itself can be performed efficiently in practice, but it sometimes outputs a large CSTP, which can impose significant overhead on the underlying SMT solver when checking its validity.) In the list-len-mode, it additionally infers invariants on the lengths of lists as described in Section 5.2 by using the backend CHC solver for integer arithmetic.
The procedure explained so far can only prove the satisfiability of the CHCs. Thus, CHoCoL also runs in parallel a procedure based on symbolic execution to prove the unsatisfiability. Given a CHC whose goal clause is $\bigwedge_{i}P_{i}(\widetilde{y}_{i})\land C\Rightarrow\mathit{false}$, CHoCoL searches for an assignment to variables $\widetilde{y}_{i}$ that makes the body ($\bigwedge_{i}P_{i}(\widetilde{y}_{i})\land C$) of the goal clause valid by recursively unfolding the predicates in the goal in a breadth-first manner. The details of this procedure are omitted, since it is rather straightforward and contains few novel ideas. It would be possible to utilize the result of tuple pattern inference also to speed-up the refutation process, but that is left for future work.
@@ 6.2 Experiments
We have conducted experiments on a benchmark set selected from CHC-COMP 2025 ADT-LIN category (https://chc-comp.github.io/). The benchmark set consists of 445 instances of the CHC satisfiability problem, which are all the instances of the ADT-LIN category where the used ADTs are only list-like data structures. The experiments were conducted on a machine with AMD Ryzen 9 5900X CPU and 32GB RAM. The cpu and wallclock time limits were 3 minutes, and the memory limit was 8GB. The number of CPU cores was limited to 4.
We compared CHoCoL with the state-of-the-art CHC solvers that support ADTs: RInGen [36], Spacer/Racer [35, 29], Eldarica [25], HoIce [7, 8], and Catalia [33]. The results are shown in Table 1 and Figure 1. To confirm that our method serves as a complement to previous methods, we have also prepared a combination of CHoCoL and Catalia, shown as CHoCoL+Catalia in the table and figure; it runs CHoCoL and Catalia in parallel. [^9] A combination of CHoCoL (without seq-cstp-mode) and Catalia, called “ChocoCatalia” won the ADT-LIN category of CHC-COMP 2025; the configuration of CHoCoL+Catalia reported in this submission slightly differs from “ChocoCatalia”.
@@
CHoCoL 180 ( 97) 80 ( 0) 260 ( 97)
RInGen 64 ( 23) 37 ( 4) 101 ( 27)
Spacer 20 ( 2) 90 ( 0) 110 ( 2)
Eldarica 20 ( 0) 84 ( 0) 104 ( 0)
HoIce 25 ( 6) 45 ( 0) 70 ( 6)
Catalia 87 ( 11) 87 ( 0) 174 ( 11)
CHoCoL+Catalia 194 84 278
@@
Table 1 shows the numbers of solved instances by CHoCoL and the other CHC solvers. For each solver, the column Solved (SAT) (resp. Solved (UNSAT)) shows the number of satisfiable (resp. unsatisfiable) instances solved, and the column Solved (all) shows the total number of instances solved. The numbers in parentheses indicate instances uniquely solved by the solver, i.e., those that no other solver (except CHoCoL+Catalia) could solve within the time limit.
The experimental results demonstrate the effectiveness of CHoCoL. CHoCoL achieved the highest total number of solved instances (260) among all compared solvers, showing its overall superiority in handling CHCs with list-like data structures. A key strength of CHoCoL is its performance on satisfiable instances, where it solved 180 instances compared to 87 by Catalia and 64 by RInGen. More importantly, CHoCoL uniquely solved 97 satisfiable instances, [^10] A increase of the time limit would not significantly change the situation, because the benchmark set contains a number of instances that require reasoning about list equalities, as in the examples in Section 1, which cannot be solved by Catalia and RInGen, which abstract lists to integers or elements of a finite domain. demonstrating the effectiveness of the STP inference approach for CHC solving.
The result of CHoCoL+Catalia, though, shows that CHoCoL also has a limitation, and that it is better to use it as a complement to other solvers. Also, in contrast with the strong performance of CHoCoL for SAT instances, CHoCoL showed slightly lower performance than other solvers (except RInGen) for UNSAT instances. This is unsurprising because our current refutation procedure is quite naive. We plan to apply STP inference also to improve the refutation procedure.
@@
Total 180 97
Table 2 shows the numbers of solved SAT instances by each mode of CHoCoL. The row “refutation” shows the number of instances found to be satisfiable only by finite unfolding. Remarkably, the list-stp-mode alone solved more SAT instances than the other solvers. As shown in the row “list-len-mode”, the combination of STPs and CHC solving over integer arithmetic was also effective. There were no uniquely solved instances for list-cstp-mode; this suggests that although CSTPs are required in theory (cf. Theorem 5.4), SPTs often suffice in practice.
We have also manually inspected some of the SAT instances that could not be solved by CHoCoL. They typically involve functions such as sort, filter, map, and fold, whose input/output relations cannot be expressed by STPs. We plan to extend our framework to support those list functions, along the line suggested in Remark 4.1.
@@ 7.1 Learning Languages from Positive Samples
The problem of learning a language from only positive samples [19] has been actively studied, and various language classes have been found to be learnable. As already mentioned, most closely related to our notion of tuple patterns is Angluin’s pattern languages [2, 1] and its variations [57, 58]. Indeed, “erasing” pattern languages [57, 58] (in which empty strings may be substituted for variables, as opposed to Angluin’s original pattern languages) can be considered singleton tuple patterns, and conversely, if we introduce a special symbol \$ (which must not occur in strings substituted for variables; such a restriction may be expressed by patterns with regular constraints [57]), then a tuple pattern $(p_{1},\ldots,p_{n})$ can be expressed as the pattern $p_{1}\mathdollar\cdots\mathdollar p_{n}$. No polynomial time learning algorithm is known for the full class of pattern languages (and in fact, the existence of such an algorithm is highly unlikely given the result of  [1]). Various subclasses of pattern languages are known to be polynomial time learnable from positive data [1, 56, 52, 57, 44]. Among others, Angluin [1] and Shinohara [58] gave polynomial time algorithms respectively for patterns consisting of a single variable and for “regular” (erasing) patterns where each variable may occur only once. Mitchell [44] extended the latter result to quasi-regular patterns, where variables must occur the same number of times. Those subclasses are not large enough to express tuple patterns like $(xy,xyx)$ (which contains two variables, each of which occurs a different number of times).
@@ 7.2 Data-Driven Approaches to Program Verification
Various data-driven approaches to program verification have recently been proposed, especially for automated discovery of program invariants. Some of them require only positive samples [55, 26, 70], some require negative samples as well [69, 68, 54], and some require even implication examples of the form $d_{1}\land\cdots\land d_{k}\Rightarrow d$, which means “if $d_{1},\ldots,d_{k}$ are positive samples, then $d$ is also positive” [18, 7, 17, 34]. The used techniques vary from algebraic ones [55, 26], SVM [69, 68], and neural networks [54, 34]. Most of those techniques have been used mainly for the discovery of invariants on integers, and it is unclear how they can be adopted to discover invariants on list-like data structures.
Among them, the techniques of Sharma et al. [55] and Ikeda et al. [26] are technically closest to ours (although they are for discovering numerical equality constraints), in that those methods can also quickly infer equality constraints only from a small number of examples. In fact, our STP inference algorithm, which repeatedly simplifies learning data in matrix form, has been inspired by those techniques. Ikeda et al. [26] applied that technique to CHC solving. The application to CHC solving is less trivial in our context, however. In the case of Ikeda et al. [26], which finds equality constraints on integers, it is sufficient to insert those invariants as a part of preprocessing, and then invoke an off-the-shelf CHC solver. For example, if we can find an invariant $P(x,y)\Rightarrow x+2y=3$, then we can provide that information to a CHC solver, just by replacing every occurrence of $P(z,w)$ in CHC with $P(z,w)\land z+2w=3$. In contrast, since there are no CHC solvers that can directly reason about sequence equality constraints, we had to develop a more sophisticated method for applying the result of CSTP inference to CHC solving, as discussed in Section 5. Actually, Theorem 5.1 can also be adopted for the method of Ikeda et al. [26] as follows: given a system $\mathcal{S}$ of CHCs over linear integer arithmetic, it is decidable whether $\mathcal{S}$ has a model that can be described solely by conjunctions of linear integer equalities of the form $\bigwedge_{i}\vec{c}_{i}\vec{x}_{i}+d_{i}=0$.
Zhu et al. [70] proposed a data-driven approach to learning invariants of (immutable) data structures. Their technique learns relations between two nodes of a data structure (e.g., “for every cons-cell $N_{1}$ in a list, the value of every cell reachable from $N_{1}$ is no less than that of $N_{1}$”), and can therefore reason about properties such as the sortedness of a list. To our knowledge, however, their technique cannot be applied to reasoning about the logical equivalence of lists; in fact, their abstraction for lists cannot distinguish [1;1;1;2;2] from [1;1;2;2;2]. Therefore, their technique cannot handle either of our motivating examples in Section 1. Related techniques have been developed for verifying programs that manipulate mutable data structures [31, 27], but these techniques also cannot reason about the logical equality of lists.
@@ 7.3 CHC Solving for Algebraic Data Types
Solving CHCs over algebraic data types (or, automated verification of programs manipulating ADTs) often require finding and reasoning about inductive properties. To this end, Unno et al. [62] introduced an automated method for applying induction in CHC solving. De Angelis et al. [13] later proposed a method to achieve similar effects without explicit induction, by using unfold/fold transformations. Mordvinov and Fedyukovich [45] also proposed a variation of unfold/fold transformations. Losekoot et al. [38, 39] applied automata-based techniques to reason about relational properties of ADTs, which also have a similar effect to unfold/fold transformation-based techniques. These methods tend to be fragile, requiring a lot of heuristics to automatically decide how and when to apply induction, unfold/fold transformations, etc.
Several methods for CHC solving have also been proposed, which abstract ADTs to elements of a finite domain or integers. Kostyukov [36] proposed a method for abstracting ADTs using automata, while the Eldarica CHC solver [25] abstracts ADTs by their size. Govind V. K. et al. [29] proposed a method that uses recursively defined functions (RDTs) over ADTs for abstraction, but those RDTs must be explicitly specified. Katsura et al. [33] proposed a method for automatically finding a catamorphism-based abstraction for ADTs. While these methods are effective for certain CHCs, they are not effective for proving the equality of ADTs; although the abstraction by using an abstraction map $\alpha$ may be used for proving $\alpha(l_{1})=\alpha(l_{2})$, it does not imply $l_{1}=l_{2}$ in general. Thus, those methods cannot be used to prove neither of the examples in Section 1. As confirmed in our experiments, our method using CSTPs is complementary to those methods; in fact, the combination of our method with Catalia achieved the best performance in the experiments in Section 6, and also in the competition CHC-COMP 2025.
@@ 8 Conclusions
We have introduced the new notions of solvable tuple patterns (STPs) and and conjunctive STPs (CSTPs), and developed their theories. In particular we have shown that both STPs and CSTPs are learnable from only positive samples, and that the satisfiability of quantifier-free formulas is decidable. We have applied those theories to prove that it is decidable whether a system of CHCs over words has a CSTP-describable model. We have implemented a new CHC solver called CHoCoL based on the proposed approach, and confirmed the effectiveness of the approach through experiments.
Future work includes: (i) generalization of STPs to tree patterns (like “f(x,g(y,x))f(x,g(y,x))”) to handle tree-like data structures, (ii) extension of STPs to infer Boolean combinations of STPs, like $(x,y,xz)\lor(x,y,yz)$ (which means “either the first or second element is a prefix of the third element), (iii) applications of other classes of languages learnable from positive samples to CHC solving, and (iv) extension and generalization of the method to combine CHC solvers over different domains.